OpenStudy (anonymous):

Solve Each Equation: First Question: 5-3(w + 4) = w-7 Second Question: 2x-5(x - 3) = 2(x-10)

OpenStudy (anonymous):

do you want the answer for w or what

OpenStudy (anonymous):

5-3w -12w=w-7 5-15w=w-7 5=16w-7 12=16w w=12/16=3/4

OpenStudy (anonymous):

2x-5x+15=2x-20 -3x+15=2x-20 -5x=-35 5x=35 x=7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np